# Thermodynamic stability in an Einstein universe

### v1

@@ 1 Introduction
Six decades ago, Penzias and Wilson found that a nearly 3K electromagnetic blackbody radiation fills up the cosmos [1]. As is well known, such a discovered have enormous implications, since e.g., it is related with the very beginning of our universe. Besides electromagnetic radiation, which is described by a massless vector field, the universe may house other types of radiation, such as scalar and spinor fields. Thus, it is certainly worth investigating the relationship between their thermodynamics and the topology and geometry of the universe.
@@ 2 Renormalized Feynman propagator
Noting eq. (16) and taking into account eqs. (11) and (12), we see that $G_{{\tt vacuum}}$ does not depend on $T$ and vanishes if $a\rightarrow\infty$, $G_{{\tt thermal}}$ vanishes if $T\rightarrow 0$, and that $G_{{\tt mixed}}$ vanishes if $a\rightarrow\infty$ or if $T\rightarrow 0$ alike. These asymptotic behaviors of the terms in eq. (16) justify their subscript.
@@ 4 $\left<T^{\mu}{}_{\nu}\right$>
$\hskip 25.6073pt=\left<T_{\mu}{}_{\nu}\right>_{{\tt vacuum}}+\left<T_{\mu}{}_{\nu}\right>_{{\tt mixed}}+\left<T_{\mu}{}_{\nu}\right>_{{\tt thermal}}$ (36)
@@ 4.1.1 $\rho_{{\tt vacuum}}$
which reproduces the well known result in eq. (5) when $\xi=1/6$ (conformal coupling), and it agrees with ref. [15] where $\zeta$-function regularization has been used. The plots in Figures 4 and 5 show a rather unsuspected behavior. Namely, although $\rho_{{\tt vacuum}}$ is positive at $\xi=1/6$ [cf. eq. (42)], as $\xi$ decreases from $1/6$, $\rho_{{\tt vacuum}}$ changes sign turning up to be negative when $\xi=0$ (minimal coupling). Such a behavior is in agreement with the study in ref. [17], but it is at variance with refs. [29, 30]  [^13] Ref. [29] claims that $\rho_{{\tt vacuum}}>0$ for $\xi\in[0,1/6]$ whereas ref. [30] states that $\rho_{{\tt vacuum}}=0$ when $\xi=0$. We have checked that these discrepancies are due to incorrectness in their calculations..
@@ 4.1.4 Conformally invariant $\rho$
which are well known expressions that have been discovery in the early papers on quantum fields in the EU (see section 1). It should be pointed out that the blackbody term in eq. (50) is corrected only by an exponentially small term containing the size of the universe. Nevertheless, we stress, this is not the case when $\xi\neq 1/6$ [cf. eq. (44)], a fact that will be relevant in the study of thermodynamic stability as has been already mentioned in the text. Perhaps it is worth remarking that eq. (49) as well as its counterpart eq. (30) still hold when $Ta\rightarrow 0$.
@@ 6 Final remarks
In this paper we have derived the renormalized Feynman propagator, at finite temperature $T$, for a massive and neutral scalar field $\phi$ in the EU of radius $a$, for arbitrary real values of the curvature coupling parameter $\xi$. A thorough analysis on how the ensemble averages $\left<\phi^{2}\right$> and $\left<T^{\mu}{}_{\nu}\right$> depend on $T$ and $\xi$ has been presented, revealing that when $Ta\rightarrow 0$ only $\xi\geq 1/6$ are allowed by thermodynamic stability of a massless $\phi$. It follows that the minimal coupling, $\xi=0$, should be discarded at low temperatures and/or small radius of the EU. Similarly, when $Ta\rightarrow\infty$, only $\xi\leq 1/6$ meets the criteria of thermodynamic stability of the massless scalar blackbody radiation in the EU. We then applied our findings to a mixture of scalar bosons, neutrinos and photons showing that at least one scalar field must be present such that the hot atmosphere is thermodynamic stable with $\xi$ tuned as in eq. (63). Therefore one concludes that neutrinos as well as photons cannot be alone in the EU when $Ta\rightarrow\infty$, which sounds like a surprise.
@@
An early paper by Hu (see ref. [31], and references therein) has explored thermodynamic equilibrium in non stationary universes  [^19] We wish to report two apparent typos in ref. [31]. To be consistent with ref. [32], the term $m^{2}T^{2}/24a^{2}$ in eq. (8) should be replaced by $m^{2}T^{2}/24$. And the factor $T^{2}/24$ in eq. (10), which is a generalization of eq. (53) in the present account, should be $T^{2}/24$ instead.. In particular for massless $\phi$ in Robertson-Walker universes, it has been shown that only $\xi=1/6$ is consistent with the notion of thermodynamic equilibrium at all times. The author then introduces the notion of quasi-thermodynamic equilibrium, according to which $\xi>1/6$ would also be allowed as long as $(6\xi-1)R/6\ll T^{2}$, where $R$ is the Ricci scalar curvature. Now, thermodynamics is also crucially about stable thermodynamic equilibrium [18]. Thus, if the material in the present paper were extended to Robertson-Walker universes, in the light of our results, most likely $\xi>1/6$ at high temperatures would be discarded by thermodynamic stability; though this point needs further investigation.
@@
We wish now to indulge in a speculation regarding the behavior of $\left<\phi^{2}\right$> and $\left<T^{\mu}{}_{\nu}\right$> when $Ta\rightarrow 0$ and $\xi<1/6$. As we know well by now, these averages will present oscillatory cusp-like patterns that could be eliminated by simply setting $\xi\geq 1/6$. Nevertheless, one might wonder if these patterns are not telling us more than that we cannot have $\xi<1/6$ near the absolute zero of temperature and/or $a\rightarrow 0$. Perhaps this point also deserves further investigation.
@@ A.1 Geometric objects
$G_{00}=-\frac{3}{a^{2}},\hskip 14.22636ptG_{11}=1,\hskip 14.22636ptG_{22}=\mathrm{sin}^{2}\chi,\hskip 14.22636ptG_{33}=\mathrm{sin}^{2}\chi\,\mathrm{sin}^{2}\theta.$
@@ A.2 Feynman propagator at finite $T$
$W_{p\,n}^{k}(\chi,\theta,\phi)=\left[\frac{2^{2n+1}(p+1)\Gamma(p-n+1)\Gamma^{2}(n+1)}{\pi\Gamma(p+n+2)}\right]^{1/2}\sin^{n}\chi C_{p-n}^{(n+1)}(\cos\chi)Y_{n}^{k}(\theta,\phi);$ (74)
@@
$\int_{0}^{\pi}d\chi\int_{0}^{\pi}d\theta\int_{0}^{2\pi}d\varphi\;W_{p\,n}^{k}(\chi,\theta,\phi)\,W_{p^{\prime}\,n^{\prime}}^{*\,k^{\prime}}(\chi,\theta,\phi)\sin^{2}\chi\sin\theta=\delta_{p\,p^{\prime}}\delta_{n\,n^{\prime}}\delta_{k\,k^{\prime}}.$ (75)
@@
$\sum_{p=0}^{\infty}\sum_{n=0}^{p}\sum_{k=-n}^{n}W_{p\,n}^{k}(\chi,\theta,\phi)\,W_{p\,n}^{*\,k}(\chi^{\prime},\theta^{\prime},\phi^{\prime})=\frac{\delta(\chi-\chi^{\prime})\delta(\theta-\theta^{\prime})\delta(\varphi-\varphi^{\prime})}{\sin^{2}\chi\sin\theta}.$ (76)
@@
$G_{E}({\rm x},{\rm x}^{\prime})=i\hskip-5.69046pt\sum_{m=-\infty}^{\infty}\sum_{p=0}^{\infty}\sum_{n=0}^{p}\sum_{k=-n}^{n}\int_{0}^{\infty}d\omega e^{-i\omega E_{m,p}}\psi_{m,p,n,k}({\rm x})\psi^{*}_{m,p,n,k}({\rm x}^{\prime}),$ (78)
@@
$G_{E}({\rm x},{\rm x}^{\prime})=-\frac{i}{\sin\alpha}\sum_{m=-\infty}^{\infty}\sum_{n=-\infty}^{\infty}(\alpha-2\pi n)\int_{0}^{\infty}\frac{d\omega}{(4\pi\omega)^{2}}e^{i\{[(\tau-\tau^{\prime}-m\beta)^{2}+(\alpha-2\pi n)^{2}a^{2}]/4\omega-\mu^{2}\omega\}}.$
@@
$G_{E}({\rm x},{\rm x}^{\prime})=\frac{\mu}{4\pi^{2}\sin\alpha}\sum_{m=-\infty}^{\infty}\sum_{n=-\infty}^{\infty}(\alpha-2\pi n)\frac{K_{1}\left(\mu\sqrt{(\tau-\tau^{\prime}-m\beta)^{2}+(\alpha-2\pi n)^{2}a^{2}}\right)}{\sqrt{(\tau-\tau^{\prime}-m\beta)^{2}+(\alpha-2\pi n)^{2}a^{2}}}.$ (83)

### v2

@@ 1 Introduction
Six decades ago, Penzias and Wilson found that a nearly 3K electromagnetic blackbody radiation fills up the cosmos [1]. As is well known, such a discovery has enormous implications, since it is related with the very beginning of our universe. Besides electromagnetic radiation, which is described by a massless vector field, the universe may house other types of radiation, such as scalar and spinor fields. Thus, it is certainly worth investigating the relationship between their thermodynamics and the topology and geometry of the universe.
@@ 2 Renormalized Feynman propagator
Noting eq. (16) and taking into account eqs. (11) and (12), we see that $G_{{\tt vacuum}}$ does not depend on $T$ and vanishes if $a\rightarrow\infty$, $G_{{\tt thermal}}$ vanishes if $T\rightarrow 0$, and that $G_{{\tt mixed}}$ vanishes if $a\rightarrow\infty$ or if $T\rightarrow 0$ alike. These asymptotic behaviors of the terms in eq. (16) justify their subscripts.
@@ 4 $\left<T^{\mu}{}_{\nu}\right$>
$\hskip 29.59067pt=\left<T_{\mu}{}_{\nu}\right>_{{\tt vacuum}}+\left<T_{\mu}{}_{\nu}\right>_{{\tt mixed}}+\left<T_{\mu}{}_{\nu}\right>_{{\tt thermal}}$ (36)
@@ 4.1.1 $\rho_{{\tt vacuum}}$
which reproduces the well known result in eq. (5) when $\xi=1/6$ (conformal coupling), and it agrees with ref. [15] where $\zeta$-function regularization has been used. The plots in Figures 4 and 5 show a rather unsuspected behavior. Namely, although massless $\rho_{{\tt vacuum}}$ is positive at $\xi=1/6$ [cf. eq. (42)], as $\xi$ decreases from $1/6$, $\rho_{{\tt vacuum}}$ changes sign turning up to be negative when $\xi=0$ (minimal coupling). Such a behavior is in agreement with the study in ref. [17], but it is at variance with refs. [29, 30]  [^13] Ref. [29] claims that $\rho_{{\tt vacuum}}>0$ for $\xi\in[0,1/6]$ whereas ref. [30] states that $\rho_{{\tt vacuum}}=0$ when $\xi=0$. .
@@ 4.1.4 Conformally invariant $\rho$
which are well known expressions that have been discovery in the early papers on quantum fields in the EU (see section 1). It should be pointed out that the blackbody term in eq. (50) is corrected only by an exponentially small term containing the size of the universe. Nevertheless, we stress, this is not the case when $\xi\neq 1/6$ where the contribution due to the finite size of the EU is much larger [cf. eq. (44)], a fact that will be relevant in the study of thermodynamic stability as has been already mentioned in the text. Perhaps it is worth remarking that eq. (49) as well as its counterpart eq. (30) still hold when $Ta\rightarrow 0$.
@@ 6 Final remarks
In this paper we have derived the renormalized Feynman propagator, at finite temperature $T$, for a massive and neutral scalar field $\phi$ in the EU of radius $a$, for arbitrary real values of the curvature coupling parameter $\xi$. A thorough analysis on how the ensemble averages $\left<\phi^{2}\right$> and $\left<T^{\mu}{}_{\nu}\right$> depend on $T$ and $\xi$ has been presented, revealing that when $Ta\rightarrow 0$ only $\xi\geq 1/6$ are allowed by thermodynamic stability of a massless $\phi$. It follows that the minimal coupling, $\xi=0$, should be discarded at low temperatures and/or small radius of the EU. Similarly, when $Ta\rightarrow\infty$, only $\xi\leq 1/6$ meets the criteria of thermodynamic stability of the massless scalar blackbody radiation in the EU. We have shown a discrepancy between statistical mechanics and the local quantum field approach for $\xi\neq 1/6$, which seems to reflect the non equivalence of global equilibrium and a local operator based notion of thermodynamic equilibrium in compact and curved spacetimes. We then applied our findings to a mixture of scalar bosons, neutrinos and photons showing that at least one scalar field must be present such that the hot atmosphere is thermodynamic stable with $\xi$ tuned as in eq. (63). Therefore one concludes that neutrinos as well as photons cannot be alone in the EU when $Ta\rightarrow\infty$, which sounds like a surprise.
@@
An early paper by Hu (see ref. [31], and references therein) has explored thermodynamic equilibrium in non stationary universes  [^19] We wish to report two apparent typos in ref. [31]. To be consistent with ref. [32], the term $m^{2}T^{2}/24a^{2}$ in eq. (8) should be replaced by $m^{2}T^{2}/24$. And the factor $T^{2}/24$ in eq. (10), which is a generalization of eq. (53) in the present account, should be $T^{2}/24$ instead. These points do not affect the conclusions in ref. [31].. In particular for massless $\phi$ in Robertson-Walker universes, it has been shown that only $\xi=1/6$ is consistent with the notion of thermodynamic equilibrium holding at all times. The author then introduced the notion of quasi-thermodynamic equilibrium, according to which $\xi>1/6$ would also be allowed as long as $(6\xi-1)R/6\ll T^{2}$, where $R$ is the Ricci scalar curvature. Now, thermodynamics is also about stable thermodynamic equilibrium [18]. Thus, it remains to be seen whether $\xi>1/6$ at high temperatures would not also be discarded by thermodynamic stability in Robertson-Walker universes. This point needs careful investigation beyond the scope of our paper.
@@
We wish now to indulge in a speculation regarding the behavior of $\left<\phi^{2}\right$> and $\left<T^{\mu}{}_{\nu}\right$> when $Ta\rightarrow 0$ and $\xi<1/6$. As we know well by now, these averages will present oscillatory cusp-like patterns that could be eliminated by simply setting $\xi\geq 1/6$, in which case their physical meaning would be just an indication of the breakdown of thermodynamic consistency. Nevertheless, one might wonder if these pathological patterns are not telling us more than that we cannot have $\xi<1/6$ near the absolute zero of temperature and/or $a\rightarrow 0$. Perhaps this point also deserves further investigation.
@@ A.1 Geometric objects
$G_{00}=-\frac{3}{a^{2}},\hskip 14.22636ptG_{11}=1,\hskip 14.22636ptG_{22}=\mathrm{sin}^{2}\chi,\hskip 14.22636ptG_{33}=\mathrm{sin}^{2}\chi\,\mathrm{sin}^{2}\theta$
@@ A.2 Feynman propagator at finite $T$
$W_{p\,n}^{k}(\chi,\theta,\phi)=\left[\frac{2^{2n+1}(p+1)\Gamma(p-n+1)\Gamma^{2}(n+1)}{\pi\Gamma(p+n+2)}\right]^{1/2}\sin^{n}\chi C_{p-n}^{(n+1)}(\cos\chi)Y_{n}^{k}(\theta,\phi)$ (74)
@@
$\int_{0}^{\pi}d\chi\int_{0}^{\pi}d\theta\int_{0}^{2\pi}d\varphi\;W_{p\,n}^{k}(\chi,\theta,\phi)\,W_{p^{\prime}\,n^{\prime}}^{*\,k^{\prime}}(\chi,\theta,\phi)\sin^{2}\chi\sin\theta=\delta_{p\,p^{\prime}}\delta_{n\,n^{\prime}}\delta_{k\,k^{\prime}}$ (75)
@@
$\sum_{p=0}^{\infty}\sum_{n=0}^{p}\sum_{k=-n}^{n}W_{p\,n}^{k}(\chi,\theta,\phi)\,W_{p\,n}^{*\,k}(\chi^{\prime},\theta^{\prime},\phi^{\prime})=\frac{\delta(\chi-\chi^{\prime})\delta(\theta-\theta^{\prime})\delta(\varphi-\varphi^{\prime})}{\sin^{2}\chi\sin\theta}$ (76)
@@
$G_{E}({\rm x},{\rm x}^{\prime})=i\hskip-5.69046pt\sum_{m=-\infty}^{\infty}\sum_{p=0}^{\infty}\sum_{n=0}^{p}\sum_{k=-n}^{n}\int_{0}^{\infty}d\omega e^{-i\omega E_{m,p}}\psi_{m,p,n,k}({\rm x})\psi^{*}_{m,p,n,k}({\rm x}^{\prime})$ (78)
@@
$G_{E}({\rm x},{\rm x}^{\prime})=-\frac{i}{\sin\alpha}\sum_{m=-\infty}^{\infty}\sum_{n=-\infty}^{\infty}(\alpha-2\pi n)\int_{0}^{\infty}\frac{d\omega}{(4\pi\omega)^{2}}e^{i\{[(\tau-\tau^{\prime}-m\beta)^{2}+(\alpha-2\pi n)^{2}a^{2}]/4\omega-\mu^{2}\omega\}}$
@@
$G_{E}({\rm x},{\rm x}^{\prime})=\frac{\mu}{4\pi^{2}\sin\alpha}\sum_{m=-\infty}^{\infty}\sum_{n=-\infty}^{\infty}(\alpha-2\pi n)\frac{K_{1}\left(\mu\sqrt{(\tau-\tau^{\prime}-m\beta)^{2}+(\alpha-2\pi n)^{2}a^{2}}\right)}{\sqrt{(\tau-\tau^{\prime}-m\beta)^{2}+(\alpha-2\pi n)^{2}a^{2}}}$ (83)